# Estimating Causal Effects from Learned Causal Networks

### v1

@@ Abstract
The standard approach to answering an identifiable causal-effect query (e.g., $P(Y|do(X)$) when given a causal diagram and observational data is to first generate an estimand, or probabilistic expression over the observable variables, which is then evaluated using the observational data. In this paper, we propose an alternative paradigm for answering causal-effect queries over discrete observable variables. We propose to instead learn the causal Bayesian network and its confounding latent variables directly from the observational data. Then, efficient probabilistic graphical model (PGM) algorithms can be applied to the learned model to answer queries. Perhaps surprisingly, we show that this model completion learning approach can be more effective than estimand approaches, particularly for larger models in which the estimand expressions become computationally difficult.
We illustrate our method’s potential using a benchmark collection of Bayesian networks and synthetically generated causal models.
[…]
1646
@@ 1 Introduction
Structural Causal Models (SCMs) are a formal framework for reasoning about causal knowledge in the presence of uncertainty [23]. When the full SCM is available, it is possible to use standard probabilistic inference [6, 8] to directly answer causal queries that evaluate how forcing some variable’s assignment $X=x$ affects another variable $Y$, written as $P(Y|do(X=x))$. However, in practice the full causal model is rarely available; instead, only a causal diagram – a directed graph capturing the causal relationships of the underlying SCM – is assumed to be known. Causal diagrams typically include both observable variables, which can be measured from data, and latent variables, which are unobservable and for which data cannot be collected. A linchpin of causal reasoning is that many causal queries can be uniquely answered using only the observable variables [28, 24, 25, 11] and consequently estimated using observational data.
@@
In this paper, we propose a straightforward yet under-explored alternative: to learn the causal model, including latent dependencies, directly from the observed data. Although the domains and distributional forms of these latent variables are unknown, recent work [34] has shown that, if the visible variables are discrete, any SCM is equivalent to one with discrete latent variables, and gives an upper bound on their domain sizes. This allows us to apply well-known techniques for learning latent variable models, such as the Expectation-Maximization (EM) algorithm, to build a Causal Bayesian Network (CBN) over the observed and latent variables. We can further apply model selection techniques such as the Bayesian Information Criterion (BIC, [16]) to select appropriate domain sizes. Then, given our learned model, we can use efficient algorithms for reasoning in probabilistic graphical models (PGMs) to answer one, or even many causal queries [6, 16].
Perhaps surprisingly, we show that this approach is often significantly more effective than the estimand-based methodology. Furthermore, we provide structural conditions that help decide when each approach is likely to be more effective. The computational benefit of our approach is tied to the complexity of the causal graph: if the causal graph has low treewidth, then exact PGM algorithms (e.g., bucket elimination or join-tree scheme) are efficiently applicable. As an added benefit if there are multiple causal queries to perform on the same model, since the model is learned only once, the learning time can be amortized over all such queries.
@@ 2 Background
We first provide some definitions and technical background. We use capital letters $X$ to represent variables, and lower case $x$ for their values. Bold uppercase ${\boldsymbol{X}}$ denotes a collection of variables, with $|{\boldsymbol{X}}|$ its cardinality and $\mathcal{D}({\boldsymbol{X}})$ their joint domain, while $\boldsymbol{x}$ indicates an assignment in that joint domain. We denote by $P({\boldsymbol{X}})$ the joint probability distribution over ${\boldsymbol{X}}$ and $P({\boldsymbol{X}=\boldsymbol{x}})$ the probability of ${\boldsymbol{X}}$ taking configuration $\boldsymbol{x}$ under this distribution, which we abbreviate $P(\boldsymbol{x})$. Similarly, $P({\boldsymbol{Y}}|{\boldsymbol{X}})$ represents the set of conditional distributions $P({\boldsymbol{Y}}\mid{\boldsymbol{X}}=\boldsymbol{x})$, $\forall\boldsymbol{x}$.
@@
An SCM $\mathcal{M}$ induces a causal diagram which is a directed graph $\mathcal{G}=\langle\boldsymbol{V}\!\cup\boldsymbol{U},E\rangle$, where each node in the graph maps to a variable in the SCM, and there is an arc from every observed or latent node $X$ in the graph to $V_{i}$ iff $X\in P\!A_{i}$. An SCM whose latent variables connect to at most a single observable variable is referred to as Markovian, and one whose latent variables connect to at most two observable variables is called Semi-Markovian. It is known that any SCM can be transformed into an equivalent Semi-Markovian one such that answers to causal queries are preserved [27]. In the semi-Markovian case it is common to use a simplified causal diagram called an Acyclic Directed Mixed Graph (ADMG), which omits any latent variables with a single child, and replaces any latent variable with two children with a bidirectional dashed arc between the children, so that latent variables are no longer explicitly shown (see, e.g., Figure 1).
@@
$P(\boldsymbol{V})=\sum_{\boldsymbol{u}}P(\boldsymbol{V},\boldsymbol{U}=\boldsymbol{u}).$ (2)
@@
An external intervention, forcing variables $\boldsymbol{X}$ to take on value $\boldsymbol{x}$, is modeled by replacing the mechanism for each $X\in\boldsymbol{X}$ with the function $X=x$. This is formally represented by the do-operator $do(\boldsymbol{X}=\boldsymbol{x})$. Thus the interventional distribution of an SCM $\mathcal{M}$ by applying $do(\boldsymbol{X})$ is,
@@
i.e., it is obtained from Eq. (1) by truncating the factors corresponding to the variables in $\boldsymbol{X}$ and setting $\boldsymbol{X}=\boldsymbol{x}$. The effect of $do(\boldsymbol{X})$ on a variable $Y$, denoted $P(Y|do(\boldsymbol{X}))$, is defined by marginalizing all the variables other than $Y$.
Causal queries.
While we normally have no access to the full structural causal model $\mathcal{M}$ and cannot evaluate these expressions directly, it is sometimes possible to evaluate the effect of an intervention given only the observational distribution $P(\boldsymbol{V})$, specifically, if the answer is unique for any full model that is consistent with the graph and $P(\boldsymbol{V})$ [23]. More formally:
Definition 2 (Identifiability).
Given a causal diagram $\mathcal{G}$, the causal effect query $P({\boldsymbol{Y}}\mid do({\boldsymbol{X}}))$ is identifiable if any two SCMs consistent with $\mathcal{G}$ that agree on the observational distribution $P(\boldsymbol{V})$ also agree on $P(\textbf{Y}\mid do(\textbf{X}))$.
Identifiability makes the causal-effect query well-posed:
Definition 3 (Causal-effect query).
Given a causal diagram $\mathcal{G}=\langle\boldsymbol{V}\!\cup\boldsymbol{U},E\rangle$, data samples from an observational distribution $P(\boldsymbol{V})$, and an identifiable query $P({\boldsymbol{Y}}\mid do({\boldsymbol{X}}))$, the task is to compute the value of $P({\boldsymbol{Y}}\mid do({\boldsymbol{X}=x}))$.
@@
The now-standard methodology for answering causal-effect queries is to break the task into two steps. The first is the identifiability step: given a causal diagram and a query $P(\boldsymbol{Y}\mid do(\boldsymbol{X}))$, determine if the query is identifiable and if so, generate an estimand, or algebraic expression in terms of the observational distribution $P(\boldsymbol{V})$ that answers the query. A complete polynomial algorithm called ID has been developed for this task [27, 24]. The second step is estimation: use samples from the observational distribution $P(\boldsymbol{V})$ to estimate the value of the estimand expression.
@@
Unfortunately, in large models the expression elements quickly become unwieldy; Example 1 involves estimating conditional distribution terms over the entire joint configuration space of the model. In terms of scalability, this presents several practical problems as model sizes increase. The required distributions have exponentially many configurations, suggesting they may require a significant amount of data to estimate accurately. Moreover, the required marginalizations are also over high-dimensional spaces, potentially also making them computationally intractable. These issues make it difficult to apply statistically sophisticated or machine learning-based estimators [12, 13, 14, 15] in such settings.
@@
These issues motivate us to explore the effectiveness of a model completion approach, in which we directly learn the Causal Bayesan Network model, including its latent variables, and then use the resulting learned model to evaluate the query. A common technique for learning latent variable models is the EM algorithm:
@@
EM is a well-studied scheme to learn Bayesian network parameters when the graph itself is known but there are some latent variables, for which data are missing [9, 16]. EM is an iterative maximum likelihood approach that alternates between an Expectation, or E-step, and a Maximization, or M-step; see [19] for details. The E-step entails inference in the Bayesian network, which can be computationally hard in general; but when the graph has bounded treewidth [^1] The treewidth is a graph parameter characterizing its sparseness; exact inference is exponential in the treewidth [8]., the E-step can be performed using join tree or bucket elimination algorithms [19, 8]. Otherwise, approximate algorithms such as belief propagation can be used [22, 20]. While EM has also been applied to structure learning [10], in our setting the structure is assumed known via the provided causal diagram.
@@
Motivated by similar model completion ideas, [7, 5] first convert the causal diagram to a circuit, exploiting the functional form of SCM mechanisms to yield compact circuits, then learn the circuit parameters via EM, and answer the causal queries. However, their work is restricted to binary-valued variables and, to the best of our knowledge, no empirical evaluation is provided. Another related line of work explores neural network approaches for causal inference [30, 31]. In that work, the authors focus on counterfactual queries and propose learning a Neural Causal Model (NCM) whose functions are neural networks, a task which requires learning the mechanisms of the underlying SCM, which is significantly more challenging and unnecessary for our task of answering causal-effect queries. Moreover, while the authors provide a theoretical analysis, the empirical validation is only performed over very small, synthetic networks. EM has also been applied to causality in [26, 33]; in the latter, the authors assume knowledge of the functional mechanisms in the SCM, with only the distributions of the discrete latent variables unknown. In contrast, our method assumes knowledge of the graph only.
@@ 3 Learning-Based Causal Inference
The approach we propose for the causal-effect task is to first learn a full CBN $\mathcal{B}=\langle\mathcal{G},\mathcal{P}\rangle$ given the causal diagram $\mathcal{G}=\langle\boldsymbol{V}\!\cup\boldsymbol{U},E\rangle$ and samples from the observational distribution $P(\boldsymbol{V})$. Then, we can answer causal-effect queries based on the truncated formula Eq. (3) using probabilistic inference over the learned CBN. However, there could be many parameterizations $\mathcal{P}$ that are consistent with the same observational distribution $P(\boldsymbol{V})$. Luckily, the identifiability property ensures that the problem remains well-posed: as long as the query is identifiable, any of these alternative parameterizations $\mathcal{P}$ will generate the same answer:
Proposition 1.
@@
We note that if the query $P({\boldsymbol{Y}}\mid do({\boldsymbol{X}}=\boldsymbol{x}))$ is not identifiable, then CBNs $\mathcal{B}=(\mathcal{G},\mathcal{P})$ with different parameters $\mathcal{P}$ inducing the same observational distribution $P(\boldsymbol{V})$ may generate different answers to the query.
Consequently, our problem reduces to the well-studied task of estimating the parameters of a Bayesian network that includes latent variables from observational data, given the network structure. A widely used algorithm for this purpose is Expectation-Maximization (EM), which aims to maximize the likelihood of the data. To apply EM, however, we also require knowledge of the latent variables’ distributional forms.
Usefully, in the case of discrete visible variables $V$, prior work has shown that any SCM can be transformed into an equivalent SCM in which the latent variables $\boldsymbol{U}$ take on discrete, finite domains [34], and provides an upper bound on the required latent domain sizes. This allows us to assume, without loss of generality, a discrete distribution for the $U_{j}$.
However, the upper bound in [34] is conservative and often very large. Moreover, such large domain sizes are sufficient but may not be required for a given SCM. In practice, we use the discrete latent domain sizes as a complexity control mechanism, to impose an additional degree of regularity on the probability distribution over the visible variables. In order to select the appropriate amount of regularization, we take the domain sizes, $\boldsymbol{k}=\{k_{U_{i}}\}$, as hyperparameters and select their values by minimizing the model’s BIC score, which penalizes models with larger domain sizes for their increased flexibility and potential to overfit [6]:
@@
where $\mathcal{D}$ are the data samples from $P(\boldsymbol{V})$, $LL_{\mathcal{B},\mathcal{D}}$ is the log likelihood of the CBN model $\mathcal{B}$ learned via EM, and $p$ is the number of free probability parameters $\boldsymbol{\theta}$ in $\mathcal{B}$.
To optimize our CBN over both the domain sizes $\boldsymbol{k}$ and probability parameters $\boldsymbol{\theta}$, we propose a practical algorithm that searches greedily in the space of $\boldsymbol{k}$ while optimizing $\boldsymbol{\theta}$ by EM. Larger domain sizes can facilitate the learning of complex models, but require more samples and more time for convergence. For this reason, we prioritize searching the model space starting from smaller domain sizes. At the same time, having domain sizes slightly too large is unlikely to hurt performance significantly, so we adopt a simple strategy of keeping the latent domain sizes equal, i.e., set all $k_{U_{i}}=k$, and gradually increase the value $k$.
Our learning strategy thus benefits from two sources of variance reduction compared to the simple plug-in estimates: first, from the graph structure itself, which imposes some regularity on $P(\boldsymbol{V})$, and second, from preferring smaller latent domain sizes $k$ when possible, which encourages learning lower-rank distributions when supported by the data.
Once a full CBN $\mathcal{B}=\langle\mathcal{G},\mathcal{P}\rangle$ is learned, any causal query $P({\boldsymbol{Y}}\mid do({\boldsymbol{X}}=\boldsymbol{x}))$ can be answered based on Eq. (3). Let $\mathcal{G}_{\overline{\boldsymbol{X}}}$ denote the graph obtained by deleting all incoming arrows to $\boldsymbol{X}$. Let $\mathcal{P}_{\boldsymbol{X}=x}=\{P(V_{i}|P\!A_{i}),P(\boldsymbol{U})\}_{V_{i}\notin\boldsymbol{X}}\cup\{P({\boldsymbol{X}=\boldsymbol{x}})=1\}$. Let $\mathcal{B}_{\boldsymbol{X}=\boldsymbol{x}}=\langle\mathcal{G}_{\overline{\boldsymbol{X}}},\mathcal{P}_{\boldsymbol{X}=\boldsymbol{x}}\rangle$ be the truncated CBN. Then we have:
Proposition 2.
$P({\boldsymbol{Y}}\mid do({\boldsymbol{X}}=\boldsymbol{x}))$ in the CBN $\mathcal{B}$ is given by $P({\boldsymbol{Y}})$ in the truncated CBN $\mathcal{B}_{\boldsymbol{X}=\boldsymbol{x}}$, i.e.,
$P_{\mathcal{B}}({\boldsymbol{Y}}\mid do({\boldsymbol{X}}=\boldsymbol{x}))=P_{B_{\boldsymbol{X}=\boldsymbol{x}}}({\boldsymbol{Y}}).$ (6)
input :  A causal diagram $\mathcal{G}=\langle\boldsymbol{U}\!\cup\!\boldsymbol{V},E\rangle$, $\boldsymbol{U}$ latent and $\boldsymbol{V}$ observables; $\mathcal{D}$ samples from $P(\boldsymbol{V})$;
Query $Q=P({\boldsymbol{Y}}\mid do({\boldsymbol{X}=\boldsymbol{x}}))$.
@@
  // $k$= latent domain size, $BIC_{\mathcal{B}}$ = $BIC$ score of $\mathcal{B}$, $\mathcal{D}$,
// $LL_{\mathcal{B}}$ is the log-likelihood of $\mathcal{B}$, $\mathcal{D}$
  1.  Initialize: $BIC_{\mathcal{B}}\leftarrow\infty$,
  2.  If $\neg\,\mathrm{identifiable}(\mathcal{G},Q)$, terminate.

### v2

@@ Abstract
The standard approach to answering an identifiable causal-effect query (e.g., $P(Y|do(X)$) given a causal diagram and observational data is to first generate an estimand, or probabilistic expression over the observable variables, which is then evaluated using the observational data. In this paper, we propose an alternative paradigm for answering causal-effect queries over discrete observable variables. We propose to instead learn the causal Bayesian network and its confounding latent variables directly from the observational data. Then, efficient probabilistic graphical model (PGM) algorithms can be applied to the learned model to answer queries. Perhaps surprisingly, we show that this model completion learning approach can be more effective than estimand approaches, particularly for larger models in which the estimand expressions become computationally difficult. We illustrate our method’s potential using a benchmark collection of Bayesian networks and synthetically generated causal models.
@@ 1 Introduction
Structural Causal Models (SCMs) are a formal framework for reasoning about causal knowledge in the presence of uncertainty [25]. When the full SCM is available, it is possible to use standard probabilistic inference [6, 9] to directly answer causal queries that evaluate how forcing some variable’s assignment $X=x$ affects another variable $Y$, written as $P(Y|do(X=x))$. However, in practice the full causal model is rarely available; instead, only a causal diagram – a directed graph capturing the causal relationships of the underlying SCM – is assumed to be known. Causal diagrams typically include both observable variables, which can be measured from data, and latent variables, which are unobservable and for which data cannot be collected. A linchpin of causal reasoning is that given a causal diagram, many causal queries can be uniquely answered using only the observable variables [32, 28, 29, 12] and consequently estimated using observational data.
@@
In this paper, we propose a straightforward yet under-explored alternative: to learn the causal model, including latent dependencies, directly from the observed data. Although the domains and distributional forms of these latent variables are unknown, recent work [38] has shown that, if the visible variables are discrete, any SCM is equivalent to one with discrete latent variables, and gives an upper bound on their domain sizes. This allows us to apply well-known techniques for learning latent variable models, such as the Expectation-Maximization (EM) algorithm, to build a Causal Bayesian Network (CBN) over the observed and latent variables. We can further apply model selection techniques such as the Bayesian Information Criterion (BIC) [18] to select appropriate domain sizes. Then, given our learned model, we can use efficient algorithms for reasoning in probabilistic graphical models (PGMs) to answer one, or even many causal queries [6, 18].
Perhaps surprisingly, we show that this approach is often significantly more effective than the estimand-based methodology. Furthermore, we provide structural conditions that help decide when each approach is likely to be more effective. The computational benefit of our approach is tied to the complexity of the causal graph: if the causal graph has low treewidth, then exact PGM algorithms (e.g., bucket elimination or join-tree scheme [8, 17]) are efficiently applicable. As an added benefit if there are multiple causal queries to perform on the same model, since the model is learned only once, the learning time can be amortized over all such queries.
@@ 2 Background
We use capital letters $X$ or $V$ to represent variables, and lower case $x$ or $v$ for their values. Bold uppercase ${\boldsymbol{X}}$ denotes a collection of variables, with $|{\boldsymbol{X}}|$ its cardinality and $\mathcal{D}({\boldsymbol{X}})$ their joint domain, while $\boldsymbol{x}$ indicates an assignment in that joint domain. We denote by $P({\boldsymbol{X}})$ the joint probability distribution over ${\boldsymbol{X}}$ and $P({\boldsymbol{X}=\boldsymbol{x}})$ the probability of ${\boldsymbol{X}}$ taking configuration $\boldsymbol{x}$ under this distribution, which we abbreviate $P(\boldsymbol{x})$. Similarly, $P({\boldsymbol{Y}}|{\boldsymbol{X}})$ represents the set of conditional distributions $P({\boldsymbol{Y}}\mid{\boldsymbol{X}}=\boldsymbol{x})$, $\forall\boldsymbol{x}$. In the definitions ahead we will use $\boldsymbol{V}$ and $\boldsymbol{U}$ to stand for arbitrary variables in models while use $X$ and $Y$ to refer specifically to variables that appear in queries.
@@
An SCM $\mathcal{M}=\langle\boldsymbol{U},\boldsymbol{V},\boldsymbol{F},P(\boldsymbol{U})\rangle$ induces a causal diagram $\mathcal{G}=\langle\boldsymbol{V}\!\cup\boldsymbol{U},E\rangle$, where each node in the graph maps to a variable in the SCM, and there is an arc from every observed or latent node $X$ in the graph to $V_{i}$ iff $X\in P\!A_{i}$. Thus $P\!A_{i}$ are the parents nodes of node $X$ in $\mathcal{G}$. An SCM whose latent variables connect to at most a single observable variable is referred to as Markovian, and one whose latent variables connect to at most two observable variables is called Semi-Markovian. It is known that any SCM can be transformed into an equivalent Semi-Markovian one such that answers to causal queries are preserved [31]. In the semi-Markovian case it is common to use a simplified causal diagram called an Acyclic Directed Mixed Graph (ADMG), which omits any latent variables with a single child, and replaces any latent variable with two children with a bidirectional dashed arc between the children, so that latent variables are no longer explicitly shown (see, e.g., Figure 1).
@@
$P(\boldsymbol{V})=\sum_{\boldsymbol{U}}P(\boldsymbol{V},\boldsymbol{U}).$ (2)
@@
An external intervention, forcing variables $\boldsymbol{X}\subseteq\boldsymbol{V}$ to take on value $\boldsymbol{x}$, is modeled by replacing the mechanism for each $X\in\boldsymbol{X}$ with the function $X=x$. This is formally represented by the do-operator $do(\boldsymbol{X}=\boldsymbol{x})$. Thus the interventional distribution of an SCM $\mathcal{M}$ when applying $do(\boldsymbol{X})$ is defined by,
@@
i.e., it is obtained from Eq. (1) by truncating from $\mathcal{M}$ the factors corresponding to the variables in $\boldsymbol{X}$ and setting $\boldsymbol{X}=\boldsymbol{x}$. The effect of $do(\boldsymbol{X})$ on a variable $Y$, denoted $P(Y|do(\boldsymbol{X}))$, is defined by marginalizing all the variables other than $Y$.
Causal-effect queries.
If we have access to the full causal model we can answer causal-effect queries from the full model using traditional probabilistic inference. However, normally we have no access to the full model $\mathcal{M}$. Yet, it was shown that assuming only a causal graph it is often possible to evaluate the effect of an intervention from the observational distribution $P(\boldsymbol{V})$ only. This occurs if the answer is unique for any full model that is consistent with the graph and $P(\boldsymbol{V})$ [25]. More formally:
Definition 2 (Identifiability, causal-effect).
Given a causal diagram $\mathcal{G}$, the causal-effect query $P({\boldsymbol{Y}}\mid do({\boldsymbol{X}}))$ is identifiable if any two SCMs having the same $\mathcal{G}$ that agree on the observational distribution $P(\boldsymbol{V})$ also agree on $P(\textbf{Y}\mid do(\textbf{X}))$.
Identifiability makes the causal-effect query well-posed.
@@
The now-standard methodology for answering causal-effect queries is to break the task into two steps. The first is the identifiability step: given a causal diagram and a query $P(\boldsymbol{Y}\mid do(\boldsymbol{X}))$, determine if the query is identifiable and if so, generate an estimand, or an algebraic expression in terms of the observational distribution $P(\boldsymbol{V})$, that answers the query. A complete polynomial algorithm called ID has been developed for this task [31, 28]. The second step is estimation: use samples from the observational distribution $P(\boldsymbol{V})$ to estimate the value of the estimand expression.
@@
Unfortunately, in large models the expression elements quickly become unwieldy; Example 1 involves estimating conditional distribution terms over the entire joint configuration space of the model. In terms of scalability, this presents several practical problems as model sizes increase. The required distributions have exponentially many configurations, suggesting they may require a significant amount of data to estimate accurately. Moreover, the required marginalizations are over high-dimensional spaces, potentially making them computationally intractable. These issues make it difficult to apply statistically sophisticated or machine learning-based estimators [13, 14, 15, 16] in such settings.
@@
Treewidth.
The treewidth of a graph (also known as induced-width) characterizes how close is the graph to a tree structure. If the treewidth of a graph is $w$, the graph can be embedded in a tree (or a hypertree) of clusters whose sizes are bounded by $w$. Exact PGM inference is exponential in the treewidth of their graphs. For more formal definition and related properties see [8, 17, 9].
The above described issues on the estimand’s computational and estimation challenges, motivate us to explore the effectiveness of a model completion approach, in which we directly learn the Causal Bayesan Network model, including its latent variables, and then use the resulting learned model to evaluate the query. A common technique for learning latent variable models is the EM algorithm.
@@
EM is a well-studied scheme to learn Bayesian network parameters when the graph itself is known but there are some latent variables, for which data are missing [10, 18]. EM is an iterative maximum likelihood approach that alternates between an Expectation, or E-step, and a Maximization, or M-step; see [21] for details. The E-step entails inference in the Bayesian network, which can be computationally hard in general; but when the graph has bounded treewidth, the E-step can be performed using join tree or bucket elimination algorithms [21, 9]. Otherwise, approximate algorithms such as belief propagation can be used [24, 22]. While EM has also been applied to structure learning [11], in our setting the structure is assumed known via the provided causal diagram.
@@
Motivated by similar model completion ideas, [7, 5] first convert the causal diagram to a circuit, exploiting the functional form of SCM mechanisms to yield compact circuits, then learn the circuit parameters via EM, and answer the causal queries. However, their work is restricted to binary-valued variables and, to the best of our knowledge, no empirical evaluation is provided. Another related line of work explores neural network approaches for causal inference [34, 35]. In that work the authors focus on counterfactual queries and propose learning a Neural Causal Model (NCM) whose functions are neural networks, a task which requires learning the mechanisms of the underlying SCM, which is significantly more challenging and unnecessary for our task of answering causal-effect queries. Moreover, while the authors provide a theoretical analysis, empirical validation is only performed over very small synthetic networks. EM has also been applied to causality in [30, 37]; in the latter, the authors assume knowledge of the functional mechanisms in the SCM, with only the distributions of the discrete latent variables unknown. In contrast, our method assumes knowledge of the graph only.
@@ 3 Learning-Based Causal Inference
The approach we propose to answer the causal-effect query is to first learn a full CBN $\mathcal{B}=\langle\mathcal{G},\mathcal{P}\rangle$ from the given causal diagram $\mathcal{G}=\langle\boldsymbol{V}\!\cup\boldsymbol{U},E\rangle$ and from samples from the observational distribution $P(\boldsymbol{V})$. Then, we can answer causal-effect queries from the full model, employing the appropriate truncation Eq. (3), via probabilistic inference algorithms over the learned CBN. However, there could be many parameterizations $\mathcal{P}$ that are consistent with the same observational distribution $P(\boldsymbol{V})$. Luckily, as long as the query is identifiable, any of these alternative parameterizations $\mathcal{P}$ will generate the same answer:
Theorem 1.
@@
Proof.
Follows directly from the definition of identifiability. ∎
Consequently, our problem reduces to the well-studied task of estimating the parameters of a Bayesian network from the network structure when given data on a subset of latent variables is not available. A widely used algorithm for this purpose is Expectation-Maximization (EM), which aims to maximize the likelihood of the data. To apply EM, however, we also require knowledge of the latent variables’ distributional forms.
Usefully, in the case of discrete visible variables $\boldsymbol{V}$, prior work has shown that any SCM can be transformed into an equivalent SCM in which the latent variables $\boldsymbol{U}$ take on discrete, finite domains [38] and it provides an upper bound on the required latent domain sizes. We will therefore assume a discrete distribution for the $U_{j}$.
However, the upper bound in [38] is conservative, and often unnecessarily large. So, we decided to use the discrete latent domain sizes as a complexity control mechanism, to impose an additional degree of regularity on the probability distribution over the visible variables. For regularization, we treat the domain sizes, $\boldsymbol{k}=\{k_{U_{i}}\}$, as hyperparameters and select their values by minimizing the model’s BIC score, which penalizes models with larger domain sizes for their increased flexibility and potential over-fitting [6]:
@@
where $\mathcal{D}$ are the data samples from $P(\boldsymbol{V})$, $LL_{\mathcal{B},\mathcal{D}}$ is the log likelihood of the CBN model $\mathcal{B}$ learned via EM, and $p$ is the number of free probability parameters, denoted $\boldsymbol{\theta}$, in the Bayesian network $\mathcal{B}$.
To optimize our CBN over both the domain sizes $\boldsymbol{k}$ and probability parameters $\boldsymbol{\theta}$, we propose a practical algorithm that searches greedily in the space of $\boldsymbol{k}$ while optimizing $\boldsymbol{\theta}$ by EM. We prioritize searching the model space starting from smaller domain sizes since simpler models are preferable (i.e., Occam Razor). We adopt a simple strategy of keeping all latent domain sizes equal, i.e., set all $k_{U_{i}}=k$, and gradually increase the value $k$.
Our learning strategy benefits from two sources of variance reduction compared to the simple plug-in estimates: first, from the graph structure, which imposes some regularity on $P(\boldsymbol{V})$; and second, by preferring smaller latent domain sizes when possible, which encourages learning lower-rank distributions when supported by the data.
Model-completion rationale.
Given any causal effect query $P({\boldsymbol{Y}}\mid do({\boldsymbol{X}}=\boldsymbol{x}))$ defined relative to causal graph $\mathcal{G}$ and given a full CBN $\mathcal{B}=\langle\mathcal{G},\mathcal{P}\rangle$ learned from observational data, Theorem 1 suggests that $P({\boldsymbol{Y}}\mid do({\boldsymbol{X}}=\boldsymbol{x}))$ can be estimated as $P({\boldsymbol{Y}})$ in the truncated learned CBN $\mathcal{B}_{\boldsymbol{X}=\boldsymbol{x}}$, i.e.,
$P({\boldsymbol{Y}}\mid do({\boldsymbol{X}}=\boldsymbol{x}))\approx P_{\mathcal{B}}({\boldsymbol{Y}}\mid do({\boldsymbol{X}}=\boldsymbol{x}))=P_{B_{\boldsymbol{X}=\boldsymbol{x}}}({\boldsymbol{Y}}).$ (6)
where $\mathcal{B}_{\boldsymbol{X}=\boldsymbol{x}}=\langle\mathcal{G}_{\overline{\boldsymbol{X}}},\mathcal{P}_{\boldsymbol{X}=\boldsymbol{x}}\rangle$ is the truncated CBN and $\mathcal{G}_{\overline{\boldsymbol{X}}}$ is the graph obtained by deleting all incoming arrows to $\boldsymbol{X}$ in $\mathcal{G}$.
input :  A causal diagram $\mathcal{G}=\langle\boldsymbol{U}\!\cup\!\boldsymbol{V},E\rangle$, $\boldsymbol{U}$ latent and $\boldsymbol{V}$ observable; samples $\mathcal{D}$ from $P(\boldsymbol{V})$;
Identifiable Query $Q=P({\boldsymbol{Y}}\mid do({\boldsymbol{X}=\boldsymbol{x}}))$.
@@
  // $k$= latent domain size, $BIC_{\mathcal{B}}\equiv BIC_{\mathcal{B},\mathcal{D}}$ the $BIC$ score of
// CBN $\mathcal{B}$ $\mathcal{D}$, $LL_{\mathcal{B}}\equiv LL_{\mathcal{B},\mathcal{D}}$ the log-likelihood of $\mathcal{B}$ $\mathcal{D}$
  1.  Initialize: $BIC_{\mathcal{B}}\leftarrow\infty$